10.A beam 5.9 m long is fixed at right end (point C) and free at the left end (point A).
A uniformly distributed load with a magnitude of 2.3 KN/m is carried by the beam
from the left to a 3.3 m span at point B. The distributed load is increased to 4.6 KN/m from point B to point C.
a. What is the magnitude of the vertical reaction (KN) at point C?
b. What is the magnitude of the moment (KN-m) at point C?
c. What is the magnitude of the shear force (KN) at point B? (indicate the sign if
negative)
d. What is the magnitude of the shear force (KN) at point C? (indicate the sign if
negative)
e. What is the magnitude of the moment (KN-m) at point B? (indicate the sign if
negative)
f. What is the magnitude of the moment (KN-m) at point C? (indicate the sign if
negative)
Answer #1

10: Given :- 2.3 kN/m 4.6 kN/m mind A K ch -3.3m. * 2.6m 5.9m Vc k 9) Magnitude of vertical reaction(kn) @ pointc; By using e2 Moment @c = {(23* 23*26423)}+{ucxaf} , (2.3*3.3) KN (4 6x2.2 kN. & 2 me # K 33 K-265 Me = {(32.25) + (31.1)} Mc. = -63.353 Magnitude of Manent ha point Bo- MB = -2.3*3.3*3.3 MB = - 12.52 KN-m 2 The moment is hogging , it is negative